OpenStudy (anonymous):

At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (raden):

let there are n people's by using combination concept : nC2 = 66 solve for n

OpenStudy (anonymous):

The first person shakes hands with n others, the second with n-1, the third with n-2 and so on until the second last shakes hands with only one other person. So we need to know the value of n such that the sum of n, n-1, n-2, n-3 etc = 66. Therefore n =11

OpenStudy (raden):

nC2 = 66 n!/(n-2)!2! = 66 n(n-1)(n-2)!/(n-2)!2! = 66 cancel out for (n-2)!, it can be n(n-1)/2 = 66 n(n-1) = 2 * 66 n(n-1) = 12 * 11 n(n-1) = 12 * (12-1) so, n = 12

OpenStudy (anonymous):

thanks